A 110-pound seafarer walks up a flight of stairs 12.8 m high in 45 seconds. How much power did he used?


1
Expert's answer
2021-05-06T07:42:09-0400

The work done by seafarer is equal to his potential energy gain:


"W = mgh"

where "m = 110lbs \approx 49.9kg" is his mass, "h = 12.8m" is the height of the stairs, and "g = 9.81N\/kg" is the gravitational acceleration.

By definition the power is the work done in unit time:


"P = \dfrac{W}{t}"

where "t = 45s" is the time.

Thus, obtain:


"P = \dfrac{mgh}{t}\\
P = \dfrac{49.9\cdot 9.81\cdot 12.8}{45} \approx 140W"

Answer. 140 W.